OpenStudy (anonymous):

In 2006, professor galen supped, from the university of missouri columbia, was awarded a presidential green challenge award for his system of converting glycerin, C3H5(OH)3 a by-product of biodiesel production, to propylene glycol, C3H6(OH)2. Propylene glycol produced in this way will be cheap enough to replace the more toxic ethylene glycol that is the primary ingredient in automobile antifreeze (A) if 50.0 mL of propylene glycol has a mass of 51.80 grams what is the density? (B) To obtain the same antifreeze protection requires 76g of propylene glycol to replace each 62g of ethylene glycol.

OpenStudy (anonymous):

Continuation in letter B: calculate the mass of propylene glycol required to replace 1.00 gal of ethylene glycol. The density of ethylene is 1.12 g/mL (C) calculate the volume of propylene glycol, in gallons needed to produce the same antifreeze protection as 1.00 gal of ethylene glycol.

OpenStudy (anonymous):

@FibonacciChick666 @watermelon14 Can i ask your help? :3 i've been solving this problem. But i can't trust my answer.

OpenStudy (watermelon14):

I am sorry I havnt learned that yet :(

OpenStudy (anonymous):

Ohh. Okaay, thanks

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (anonymous):

@robtobey

OpenStudy (theeric):

For part a, you'll want to used the density formula again. The variable for density is usually \(\rho\), so I'll use that. Mass is \(m\) and volume is \(V\). Then \(\rho=m\div V\). \(51.80\ [g]\) is your mass. \(50.0\ [mL]\) is your volume.

OpenStudy (anonymous):

Yeah, i got the answer in letter A 1.02 g/mL. Right? :3

OpenStudy (theeric):

I got \(1.036\ [g/mL]\). That would round to \(1.04\ [g/mL]\), to keep with your significant figures. Would you agree?

OpenStudy (theeric):

For part b, I don't see the question.

OpenStudy (theeric):

Nevermind! :P

OpenStudy (anonymous):

No, if you divide 51.08 g and 50.0 mL. The answer would be 1.0216. Right? so if you round it off. It would be 1.02, right? Right? XD

OpenStudy (anonymous):

Ohh, I'm sorry. You were right. XD i used 51.08 instead of 51.80. I'm sorry.

OpenStudy (theeric):

I figured it was calculator error for one of us! I'm making a table of the variables - to get organized. I'll post it in a moment!

OpenStudy (theeric):

\[\begin{matrix}\text{quantity}&\text{eth. gly.}&\text{prop. gly.}\\\rho\ [g/mL]&1.12&1.04\\V\ [gal]&1.00&?\end{matrix}\]And \(76\ [g]\text{prop.}\equiv62\ [g]\text{eth.}\) in terms of antifreeze protection..

OpenStudy (anonymous):

what letter are you answering? :)

OpenStudy (theeric):

B. And you don't know either masses. You don't know how much mass 1 gallon of ethylene glycol has. And you don't know how much mass of propylene glycol you'll need to have the same antifreeze effect. \[\rho=m\div V\implies m=\rho\ V\]1 gallon is 3,785.41 milliliter, according to Google. So we'll use that as our volume.\[m=1.12\ [g/mL]\times 3,785.41\ [mL]=4,239.6592\ [g]\approx 4.24\times 10^4\ [g]\]

OpenStudy (theeric):

You already have A.

OpenStudy (theeric):

\[\begin{matrix}\text{quantity}&\text{eth. gly.}&\text{prop. gly.}\\\rho\ [g/mL]&1.12&1.04\\V\ [gal]&1.00&?\\m\ [g]&4.24\times 10^4&?\end{matrix}\]So, we know everything about ethylene glycol. And we can find out how many grams of propylene glycol we'll need to do the same antifreeze work.\[76\ [g\ prop]=62\ [g\ eth.]\]We have the grams of ethylene glycol, and we want the grams of propylene glycol. Any idea? It's much like a conversion.

OpenStudy (anonymous):

what letter is that?

OpenStudy (theeric):

\(m\) for mass.

OpenStudy (theeric):

you mean the grams of stuff? Or which letter of the problem?

OpenStudy (anonymous):

which letter of the problem. :)

OpenStudy (theeric):

Oh, B! Let me know when you're caught up!

OpenStudy (anonymous):

ok. :) i will.

OpenStudy (theeric):

A takes a little work, B takes a bit more work, and then C will be about as bad as A. :)

OpenStudy (anonymous):

Haha.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

4.08x10^4g. I'm not pretty sure :/ @theEric

OpenStudy (anonymous):

that's the volume of propylene glycol.

OpenStudy (theeric):

Let me check!

OpenStudy (anonymous):

okay :)

OpenStudy (theeric):

Grams is a unit of mass. Did you mean that's the mass of propylene glycol?

OpenStudy (anonymous):

that's gal. not g, i was wrong. Sorry

OpenStudy (theeric):

I didn't get there yet. Part B asks for mass. What did you get for that?

OpenStudy (anonymous):

ohh wait. I will calculate it.

OpenStudy (anonymous):

69.4 mL that's my answer for the mass of propylene glycol.

OpenStudy (anonymous):

@theEric

OpenStudy (theeric):

\[76\ [g\ prop]=62\ [g\ eth.]\]\[m_{eth}=4.24\times 10^4\ [g eth]\] So convert the grams of ethylene glycol to grams of propylene glycol - just like they're units of antifreeze ability! Do you know what I mean?

OpenStudy (theeric):

Hmmm?

OpenStudy (anonymous):

Okay. I surrender, i can't get it. T_T

OpenStudy (anonymous):

I can't get the answer.

OpenStudy (theeric):

Haha, all you had to say is that you didn't know what I mean. And I'll show you.

OpenStudy (anonymous):

Haha. Stop reading my mind. XD yeaah, i admit it. :P

OpenStudy (theeric):

\[4.24\times 10^3\ [g\ eth]\times \frac{76\ [g\ prop]}{62\ [g\ eth]}\\=5.1974193548387096774193548387097\times 10^3\ [g\ prop]\\\approx 5.20\times10^3\ [g\ prop]\]

OpenStudy (theeric):

That would be B.

OpenStudy (anonymous):

Ahhh! I get it :)

OpenStudy (theeric):

Next\[\begin{matrix}\text{quantity}&\text{eth. gly.}&\text{prop. gly.}\\\rho\ [g/mL]&1.12&1.04\\V\ [gal]&1.00&?\\m\ [g]&4.24\times 10^3&5.20\times 10^3\end{matrix}\] 1 unknown left. The volume of propylene glycol. Use \(\rho=m\div V\implies V=m\div\rho\). You have \(m\) in [g] and \(\rho\) in g/mL, so you'll get your \(V\) in mL. But then you can convert, if you want.

OpenStudy (theeric):

Do you see what to do there?

OpenStudy (anonymous):

yeah. :) wait. I'll try to answer it.

OpenStudy (theeric):

Okay! I'll be here!

OpenStudy (anonymous):

5000 gal.

OpenStudy (theeric):

That doesn't seem right... Seems a little too high.

OpenStudy (theeric):

\(V=m\div \rho\) \(m=5.2\times 10^3\ [g]\) \(\rho=1.04\ [g/mL]\) Plug it into a calculator and...

OpenStudy (anonymous):

it's 5000 gal, the same with my answer.

OpenStudy (anonymous):

i mean 5000 mL

OpenStudy (anonymous):

wait, i'm wrong. I'll convert it to gallons.

OpenStudy (anonymous):

I got it. @theEric 1.32 gal

OpenStudy (anonymous):

@theEric are you still there? :3

OpenStudy (theeric):

Yep! I agree with 5000mL! I'm working on the conversion! My browser is being slow.

OpenStudy (theeric):

1.32 gallons seems reasonable, but I want to double-check the calculation for you.

OpenStudy (theeric):

I got the same as you, and that's C! :)

OpenStudy (anonymous):

Yay! :D thank you so much eric. ^_^

OpenStudy (theeric):

You're very welcome. I hope it's all right! :)